OpenStudy (anonymous):

If the equation 4x^2 + 6x + c = 0 has exactly one root, what is the value of c?

OpenStudy (anonymous):

determine c so it is a perfect square \[(x+n)^2= x^2+2nx+n^2\]

OpenStudy (anonymous):

sorry it would be \[(ax+b)^2 = (ax)^2+ 2abx+b^2\]